# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ 3.2 Numerical results for $\mathrm{COV}-\mathrm{COV}_{4}$
When $k>2$, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ cannot be easily derived explicitly and numerical computations are necessary. By using the expressions derived in Appendix A.2, alongside numerical computations for the inverse of $\mathrm{COV}$, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ can be calculated. These calculations are performed for 20 different group centers and 18 different scenarios of group proportions (see Appendices A.2 and B for details). Boxplots in Figure 2 display the outcomes when $p=q=k-1$, for $k=2$, 3, 5 and 10 groups (as in Figure 1). The boxplots consist of a single line for each scenario, indicating that varying the group means does not impact the eigenvalues when $q=k-1$. However, the values of the eigenvalues differ from one scenario to another. Therefore, it can be concluded that the group proportions do have an impact on the eigenvalues. A more detailed examination of this impact will be presented in the next section. When $q<k-1$, the results are different, as illustrated in Figure 3 for $k=3$ and $q=1$, $k=5$ and $q=1$, $k=3$ and $q=2$, and $k=3$ and $q=3$ (see also Figure 10 in Appendix B with $k=10$ and $q=1,3,5,7$). For each scenario, the eigenvalues exhibit quite large variability across the different group centers. This variability is such that boxplots cross red horizontal dotted line corresponding to the value one. Unlike the case where $q=k-1$, we cannot draw any conclusions about the conditions which cause the eigenvalues to be larger or smaller than one. The influence of group proportions on eigenvalues persists, but the influence of the group centers is also crucial.
Fig. 2: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $p=q=k-1$ and with no within-group variability, when the group centers vary across 20 different configurations (see Appendix B for details). The values of $k$ vary across the panels and the $x$-axis of each panel corresponds to 18 group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$).
Fig. 3: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $p=q<k-1$ and with no within-group variability, when the group centers vary across 20 different configurations (see Appendix B for details). The values of $k$ and $q$ vary across the panels and the $x$-axis of each panel corresponds to 18 group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$).
@@
As stated in Section 3, in the context of a mixture of Dirac distributions with $q=k-1$, only the group proportions influence the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. In this section, the group means are therefore fixed to the first row of the group centers of Table 1 of Appendix B. The relationship between the group proportions and the eigenvalues is further investigated in the same context as in the previous subsection. Theoretical calculations do not permit to measure precisely the impact of group proportions on the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. Therefore, we vary the group proportions on a grid and calculate numerically the corresponding eigenvalues to better understand their behavior. The three-group case is described in Subsection 4.1, while Subsection 4.2 generalises to $k$ groups.
@@ 4.1 The case of three groups
One of the most advantageous aspects of studying three groups is the ability to represent the proportions of each group in a ternary diagram, as in Figure 4. This diagram (see Pawlowsky-Glahn et al., 2015) displays the values of three positive variables that sum up to 100%, and can be applied for the three-group proportions. Each point on the ternary diagram represents a scenario of group proportions, and the color of the point is determined by the eigenvalues obtained for that scenario. As $k=3$, we have $q=2$ and we focus on the case $p=2$ in order to avoid irrelevant dimensions associated with eigenvalues equal to one. $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ has thus two eigenvalues $\rho_{1}\geq\rho_{2}$ and two ternary diagrams are plotted: Figure 4(a) for $\rho_{1}$, and Figure 4(b) for $\rho_{2}$.
@@
Fig. 4: Ternary diagram of the group proportions ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) color-coded by the values of $\rho_{1}$ (a) and $\rho_{2}$ (b) when $k=3$ and with no within-group variability. The color gradient indicates the departure of the eigenvalues from one (white when equal, blue when smaller and red when larger). Group means are fixed to $t_{1}=(200,0)$, $t_{2}=(400,100)$, $t_{3}=(0,0)$ but have no impact on the results. See Appendix C.1 for details.
@@
Figure 4(a) illustrates that the values of $\rho_{1}$ are less than one in the center of the plot, indicating that when none of the three groups has a proportion smaller than roughly 18%, $\rho_{1}$ is less than one. Around this region, there is a white area that corresponds (given the continuity of the eigenvalues as functions of the group proportions) to $\rho_{1}$ almost equal or equal to one. $\rho_{1}=1$ is thus indistinguishable from the eigenvalues associated with the $p-q$ dimensions that do not span the FDS. It occurs when one group proportion is roughly 18%. The area beyond this white zone in the direction of the vertices is red, which indicates that the value of $\rho_{1}$ is greater than one when a group proportion is smaller than 18%. Looking at the areas around the vertices on Figure 4(b) demonstrates that the value of $\rho_{2}$ is greater than one for scenarios where two proportions are smaller than roughly 18%. A white area is also present around each of these regions, which corresponds to values of $\rho_{2}$ almost equal or equal to one. It occurs once again when one group has a proportion of roughly 18% and another group has a proportion smaller than or equal to 18%. The rest of the plot is blue, implying $\rho_{2}$ is smaller than one. A comparison of Figures 4(a) and 4(b) indicates that the white regions do not seem to intersect, thereby suggesting that the two eigenvalues are not equal to one simultaneously. In the two-group case, at the threshold value of 21%, all eigenvalues are equal to one. In the three-group case with $q=k-1$, our conjecture is that there is at least one eigenvalue that differs from one for any possible group proportion scenario.
@@
The eigenvalues representation from Figures 4(a) and 4(b) are interesting since the greater the deviation from one, the easier it will be to detect the directions that span the FDS when in practice we will have $p>k$. To facilitate the comparison between the behavior of the two eigenvalues, both are plotted on the same ternary diagram on Figure 5.
@@
Fig. 5: Ternary diagram of the group proportions ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) color-coded by the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $k=3$ and $p=q=2$ (blue when both eigenvalues are smaller than one, purple when one eigenvalue is larger and the other is smaller than one, and red when both eigenvalues are larger than one). Group centers are fixed to $t_{1}=(200,0)$, $t_{2}=(400,100)$, and $t_{3}=(0,0)$ but have no impact on the results. See Appendix C.1 for more details.
@@
In the event that two groups exhibit proportions roughly smaller than 18%, both eigenvalues are greater than one, as may be observed in the red area of Figure 5. When a single group proportion is roughly smaller than 18%, the first eigenvalue is greater than one and the second is less than one. This occurs in the purple region in Figure 5. Finally, when all groups are in proportion roughly larger than 18%, in the blue area of Figure 5, both eigenvalues are less than one. As illustrated in Figures 4(a) and 4(b), each boundary between the colored zones should display scenarios for which a single eigenvalue differs from one. However, the analysis has a numerical constraint: the 0.1% step size for the discretization of the group proportions while we use the exact value of one as the reference eigenvalue. A consequence of this experiment limitation is that only three scenarios exhibit an eigenvalue precisely equal to one. They correspond to a proportion of 60% for one group and 20% for the other two groups. These points are not represented in Figure 5.
@@
The findings for $k=3$ indicate that there is a threshold value (around 18%) such that when a group proportion goes from a value above this threshold to a value below, an eigenvalue of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ goes from a value less than one to a value greater than one. For two groups, this threshold is known and is approximately 21%. From Figure 5, establishing the precise threshold for three groups is not easy, but it is clear that the threshold is less than 21% and approximately 18%. Additionally, the rounded boundaries between the zones suggest that this threshold may slightly vary with the other group proportions. The goal now is to increase the number of groups beyond three, and to study whether there also exists a threshold value for the group proportions that would induce the same impact on the behavior of the ICS eigenvalues with $\mathrm{COV}-\mathrm{COV}_{4}$. Subsection 4.2 also proposes to numerically approximate the values of these thresholds as a function of the number of groups $k$.
@@ 4.2 The general case of $k$ groups
The thresholds described previously are computed in this section for three setups of group proportions and for values of $k$ between 2 and 10. Note that for a given $k$, the maximum proportion of a group is larger than $1/k$, while the minimum proportion is smaller than $1/k$. The balanced case corresponds to proportions all equal to $1/k$. In Setup 1, we consider scenarios of the mixture proportions such that the first group proportion ranges from 0.001 to $1/k$, the intermediate group proportions are set to $1/k$, and the last group proportion is adjusted to maintain the total sum to one. The eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are calculated for each scenario. The threshold is defined as the first value of the first group proportion for which an eigenvalue exceeds one when increasing the first group proportion. In Setup 2, the second group proportion is set to the threshold value identified in Setup 1 plus 2%. The first group proportion ranges from 0.001 to $1/k$. The remaining groups maintain a proportion of $1/k$, with the exception of the last group, whose proportion is adjusted as in Setup 1 to maintain the total sum to one. The objective of this setup is to examine the influence of a group located in proximity to the threshold. It requires at least three groups. The threshold is determined similarly to Setup 1. The grid in Setup 3 is identical to the one in Setup 2, with the exception that the second group proportion is fixed at 5%, and the last group proportion is adjusted accordingly. This setup includes a group proportion initially below the threshold for any value of $k$ between 3 and 10, thereby ensuring that at least one eigenvalue exceeds one. The threshold is defined as the first value of the first group proportion for which two eigenvalues exceed one. As with Setup 2, this method requires at least three groups. Further details and examples may be found in Appendix C.2. The thresholds identified through the three setups, for $k=2$ to 10 groups, are plotted in Figure 6.
Fig. 6: Thresholds computed in Setups 1, 2 and 3, for different values of $k$.
@@
Since the grid generation in Setups 2 and 3 requires more than three groups, there is only one point when $k=2$ for Setup 1. This point is roughly equal to 21% which is the result already known theoretically. For $k=3$, the threshold value found for Setup 1 is around 18%, the one for Setup 2 is 20%, and the one for Setup 3 is 17%. Fixing a group proportion near the first threshold or below the threshold has indeed an impact. Nonetheless, the thresholds remain around 18%. As $k$ increases, the thresholds for the three setups appear to converge toward the same value and to depend solely on $k$. Table 5 in Appendix C.2 gives the specific values of the thresholds for $k\in\{3,\ldots,10\}$. It is worth noting that the presented setups could be applied to a larger number of groups, although we have limited our analysis to 10 groups.
@@
Once the thresholds for three to ten groups have been identified, the boxplots from the previous section can be reproduced for adjusted scenarios. It is of particular interest to visualize the eigenvalues when one group proportion is at the identified threshold, and to select two or three of the previous mixture proportion scenarios to vary the number of groups with proportions below this threshold. Figure 7 illustrates this experiment. Although boxplots are plotted, the group centers remain fixed and equal to the values of the first row of the group centers of Table 1 of Appendix B.
@@
Fig. 7: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ with no within-group variability, with $p=q=k-1$, and 14 different group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$) on the $x$-axis. The values of $k$ vary across panels. In each panel, the second scenario (on the $x$-axis) contains a group at the identified threshold value (see Appendices B and C.2 for details).
@@
The results presented in Figure 7 validate the thresholds of Setup 1 presented in Figure 6 for $k=$2, 3, 5, and 10 groups. In each boxplot, the scenario where a group is at the threshold level has an eigenvalue almost equal to one. This is the first eigenvalue, given that these scenarios do not have proportions below the respective threshold. The results are consistent with the logic described in Figure 5, which states that the number of eigenvalues greater than one is linked to the number of groups with proportion below the threshold. For the scenarios of Figure 7, when $k=5$ (resp. $k=10$), the number of eigenvalues larger than one is equal to the number of groups with proportions equal to 10% (resp. 5%). However, this relationship needs further investigation in cases involving more than three groups.
@@
In Section 3, Figures 2 and 3 display the eigenvalues of ICS using $\mathrm{COV}-\mathrm{COV}_{4}$ for various Gaussian mixture models described by equation (5). Figure 2 illustrates that, if $q=k-1$, these eigenvalues do not depend on the group means, and Proposition 1 in Section 4 confirms the result theoretically. However, as shown in Figure 3, this result does not hold if $q<k-1$, leading to a more complex analysis of the eigenvalues. In the present section, we examine the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for model (2) with three Gaussian groups ($k=3$) and $q=1<k-1=2$. This corresponds to the following $p$-dimensional Gaussian mixture:
$\alpha_{1}{\cal N}_{p}(t_{1},I_{p})+\alpha_{2}{\cal N}_{p}(t_{2},I_{p})+(1-\alpha_{1}-\alpha_{2}){\cal N}_{p}(0_{p},I_{p}),$ (6)
with $t_{1}=(t_{11},0,\ldots,0)^{\top}$, $t_{2}=(t_{21},0,\ldots,0)^{\top}$, $t_{11}\neq 0$, $t_{21}\neq 0$, and $t_{11}\neq t_{21}$. We can prove the following proposition.
@@
Let us consider model (6) and let $r_{\alpha_{1},\alpha_{2}}$ be the following polynomial of degree 4:
$r_{\alpha_{1},\alpha_{2}}(x)=\alpha_{1}(-1+7\alpha_{1}-12\alpha_{1}^{2}+6\alpha_{1}^{3})x^{4}+4\alpha_{1}\alpha_{2}(1-6\alpha_{1}+6\alpha_{1}^{2})x^{3}+6\alpha_{1}\alpha_{2}(1-2\alpha_{2}+$
$\alpha_{1}(-2+6\alpha_{2}))x^{2}+4\alpha_{1}\alpha_{2}(1-6\alpha_{2}+6\alpha_{2}^{2})x+\alpha_{2}(-1+7\alpha_{2}-12\alpha_{2}^{2}+6\alpha_{2}^{3}).$
@@
The proof is given in Appendix D. Proposition 2 demonstrates that the behavior of the ICS eigenvalues for a mixture of three Gaussian groups with aligned group centers differs from the two-group case (see Subsection 2.3). For model (6), the condition under which all eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ equal one depends not only on the group proportions but also on the group centers, as determined by the fourth-order equation (7). This equation has four solutions in the complex plane but, between zero and four solutions in the real space, depending on the group proportions. Using Mathematica (Wolfram Research, Inc., 2022) to simplify the computations, we found that for $\alpha_{1}=\alpha_{2}\leq(3-\sqrt{3})/12$ ($\simeq 10.6\%$), or $\alpha_{1}=\alpha_{2}$ equal to $1/3$ or $1/4$, equation (7) has no real solution. This implies that one and only one eigenvalue of ICS differs from one, and is associated with an eigenvector that spans the FDS (which is only one-dimensional for model (6)). However, we can also identify group proportions for which, given specific group centers, all eigenvalues of ICS equal one, indicating that ICS does not work in all situations. For instance, when $\alpha_{1}=\alpha_{2}=1/6$, equation (7) has three real solutions, and all the eigenvalues of ICS equal one if $t_{21}/t_{11}$ equal -1, $(2\sqrt{6}+7)/5$ or $(-2\sqrt{6}+7)/5$. While these situations are highly specific, they illustrate that analysing the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ becomes more complex when $q$ is less than $k-1$, and that this phenomenon is already true for $k=3$.

### v2

@@ 3.2 Numerical results for $\mathrm{COV}-\mathrm{COV}_{4}$
When $k>2$, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ cannot be easily derived explicitly and numerical computations are necessary. By using the expressions derived in Appendix A.2, alongside numerical computations for the inverse of $\mathrm{COV}$, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ can be calculated. These calculations are performed for the 20 specified group centers and across 18 different scenarios of group proportions (see Appendix A.2 and Supplementary Section S2 for details). Boxplots in Fig. 2 display the outcomes when $p=q=k-1$, for $k\in\{2,3,5,10\}$ groups (as in Fig. 1). The boxplots consist of a single line for each scenario, indicating that varying the group means does not impact the eigenvalues when $q=k-1$. However, the values of the eigenvalues differ from one scenario to another. Therefore, it can be concluded that the group proportions do have an impact on the eigenvalues. A more detailed examination of this impact will be presented in the next section. When $q<k-1$, the results are different, as illustrated in Fig. 3 for $k=3$ and $q=1$, and $k=5$ and $q\in\{1,2,3\}$ (see also Fig. S1 in Supplementary Section S2 with $k=10$ and $q\in\{1,3,5,7\}$). For each scenario, the eigenvalues exhibit quite large variability across the different group centers. This variability is such that boxplots cross red horizontal dotted line corresponding to the value one. Unlike the case where $q=k-1$, we cannot draw any conclusions about the conditions which cause the eigenvalues to be larger or smaller than one. The influence of group proportions on eigenvalues persists, but the influence of the group centers is also crucial.
Fig. 2: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $p=q=k-1$ and with no within-group variability, when the group centers vary across 20 different configurations. The values of $k$ vary across the panels and the $x$-axis of each panel corresponds to 18 group proportions scenarios ($\alpha_{j}$$j\in\{1,\ldots,k\}$).
Fig. 3: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $p=q<k-1$ and with no within-group variability, when the group centers vary across 20 different configurations. The values of $k$ and $q$ vary across the panels and the $x$-axis of each panel corresponds to 18 group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$).
@@
As stated in Section 3, in the context of a mixture of Dirac distributions with $q=k-1$, only the group proportions influence the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. In this section, the group means are therefore fixed to the first row of the group centers of Table S1 of Supplementary Section S1. The relationship between the group proportions and the eigenvalues is further investigated in the same context as in the previous subsection. Theoretical calculations do not permit to measure precisely the impact of group proportions on the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. Therefore, we vary the group proportions on a grid and calculate numerically the corresponding eigenvalues to better understand their behavior. The three-group case is described in Subsection 4.1, while Subsection 4.2 generalises to $k$ groups.
@@ 4.1 The case of three groups
One of the most advantageous aspects of studying three groups is the ability to represent the proportions of each group in a ternary diagram, as in Fig. 4. This diagram (see Pawlowsky-Glahn et al., 2015) displays the values of three positive variables that sum up to 100%, and can be applied for the three-group proportions. Each point on the ternary diagram represents a scenario of group proportions, and the color of the point is determined by the eigenvalues obtained for that scenario. As $k=3$, we have $q=2$ and we focus on the case $p=2$ in order to avoid irrelevant dimensions associated with eigenvalues equal to one. $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ has thus two eigenvalues $\rho_{1}\geq\rho_{2}$ and two ternary diagrams are plotted: Fig. 4(a) for $\rho_{1}$, and Fig. 4(b) for $\rho_{2}$.
@@
Fig. 4: Ternary diagram of the group proportions ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) color-coded by the values of $\rho_{1}$ (a) and $\rho_{2}$ (b) when $k=3$ and with no within-group variability. The color gradient indicates the departure of the eigenvalues from one (white when equal, blue when smaller and red when larger). Group means are fixed to $\mathbf{t}_{1}=(200,0)$, $\mathbf{t}_{2}=(400,100)$, $\mathbf{t}_{3}=(0,0)$ but have no impact on the results.
@@
Fig. 4(a) illustrates that the values of $\rho_{1}$ are less than one in the center of the plot, indicating that when none of the three groups has a proportion smaller than roughly 18%, $\rho_{1}$ is less than one. Around this region, there is a white area that corresponds (given the continuity of the eigenvalues as functions of the group proportions) to $\rho_{1}$ almost equal or equal to one. $\rho_{1}=1$ is thus indistinguishable from the eigenvalues associated with the $p-q$ dimensions that do not span the FDS. It occurs when one group proportion is roughly 18%. The area beyond this white zone in the direction of the vertices is red, which indicates that the value of $\rho_{1}$ is greater than one when a group proportion is smaller than 18%. Looking at the areas around the vertices on Fig. 4(b) demonstrates that the value of $\rho_{2}$ is greater than one for scenarios where two proportions are smaller than roughly 18%. A white area is also present around each of these regions, which corresponds to values of $\rho_{2}$ almost equal or equal to one. It occurs once again when one group has a proportion of roughly 18% and another group has a proportion smaller than or equal to 18%. The rest of the plot is blue, implying $\rho_{2}$ is smaller than one. A comparison of Fig. 4(a) and 4(b) indicates that the white regions do not seem to intersect, thereby suggesting that the two eigenvalues are not equal to one simultaneously. In the two-group case, at the threshold value of 21%, all eigenvalues are equal to one. In the three-group case with $q=k-1$, our conjecture is that there is at least one eigenvalue that differs from one for any possible group proportion scenario.
@@
The eigenvalues representation from Fig. 4(a) and 4(b) are interesting since the greater the deviation from one, the easier it will be to detect the directions that span the FDS when in practice we will have $p>k$. To facilitate the comparison between the behavior of the two eigenvalues, both are plotted on the same ternary diagram on Fig. 5.
@@
Fig. 5: Ternary diagram of the group proportions ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) color-coded by the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for $k=3$ and $p=q=2$ (blue when both eigenvalues are smaller than one, purple when one eigenvalue is larger and the other is smaller than one, and red when both eigenvalues are larger than one). Group centers are fixed to $\mathbf{t}_{1}=(200,0)$, $\mathbf{t}_{2}=(400,100)$, and $\mathbf{t}_{3}=(0,0)$ but have no impact on the results.
@@
In the event that two groups exhibit proportions roughly smaller than 18%, both eigenvalues are greater than one, as may be observed in the red area of Fig. 5. When a single group proportion is roughly smaller than 18%, the first eigenvalue is greater than one and the second is less than one. This occurs in the purple region in Fig. 5. Finally, when all groups are in proportion roughly larger than 18%, in the blue area of Fig. 5, both eigenvalues are less than one. As illustrated in Fig. 4(a) and 4(b), each boundary between the colored zones should display scenarios for which a single eigenvalue differs from one. However, the analysis has a numerical constraint: the 0.1% step size for the discretization of the group proportions while we use the exact value of one as the reference eigenvalue. A consequence of this experiment limitation is that only three scenarios exhibit an eigenvalue precisely equal to one. They correspond to a proportion of 60% for one group and 20% for the other two groups. These points are not represented in Fig. 5. Refer to Appendix B.1 for additional details on the construction of Fig. 4 and 5.
@@
The findings for $k=3$ indicate that there is a threshold value (around 18%) such that when a group proportion goes from a value above this threshold to a value below, an eigenvalue of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ goes from a value less than one to a value greater than one. For two groups, this threshold is known and is approximately 21%. From Fig. 5, establishing the precise threshold for three groups is not easy, but it is clear that the threshold is less than 21% and approximately 18%. Additionally, the rounded boundaries between the zones suggest that this threshold may slightly vary with the other group proportions. The goal now is to increase the number of groups beyond three, and to study whether there also exists a threshold value for the group proportions that would induce the same impact on the behavior of the ICS eigenvalues with $\mathrm{COV}-\mathrm{COV}_{4}$. Subsection 4.2 also proposes to numerically approximate the values of these thresholds as a function of the number of groups $k$.
@@ 4.2 The general case of $k$ groups
The thresholds described previously are computed in this section for three setups of group proportions and for values of $k$ between 2 and 10. Note that for a given $k$, the maximum proportion of a group is larger than $1/k$, while the minimum proportion is smaller than $1/k$. The balanced case corresponds to proportions all equal to $1/k$. In Setup 1, we consider scenarios of the mixture proportions such that the first group proportion ranges from 0.001 to $1/k$, the intermediate group proportions are set to $1/k$, and the last group proportion is adjusted to maintain the total sum to one. The eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are calculated for each scenario. The threshold is defined as the first value of the first group proportion for which an eigenvalue exceeds one when decreasing the first group proportion. In Setup 2, the second group proportion is set to the threshold value identified in Setup 1 plus 2%. The first group proportion ranges from 0.001 to $1/k$. The remaining groups maintain a proportion of $1/k$, with the exception of the last group, whose proportion is adjusted as in Setup 1 to maintain the total sum to one. The objective of this setup is to examine the influence of a group located in proximity to the threshold. It requires at least three groups. The threshold is determined similarly to Setup 1. The grid in Setup 3 is identical to the one in Setup 2, with the exception that the second group proportion is fixed at 5%, and the last group proportion is adjusted accordingly. This setup includes a group proportion initially below the threshold for any value of $k$ between 3 and 10, thereby ensuring that at least one eigenvalue exceeds one. The threshold is defined as the first value of the first group proportion for which two eigenvalues exceed one. As with Setup 2, this method requires at least three groups. Further details and examples may be found in Appendix B.2. The thresholds identified through the three setups, for $k=2$ to 10 groups, are plotted in Fig. 6.
Fig. 6: Group proportion thresholds computed for different values of the number of groups $k$ and three setups. Setup 1: all group proportions initially equal. Setup 2: $\alpha_{2}$ near the Setup 1 threshold. Setup 3: $\alpha_{2}$ equal to 5%.
@@
Since the grid generation in Setups 2 and 3 requires more than three groups, there is only one point when $k=2$ for Setup 1. This point is roughly equal to 21% which is the result already known theoretically. For $k=3$, the threshold value found for Setup 1 is around 18%, the one for Setup 2 is 20%, and the one for Setup 3 is 17%. Fixing a group proportion near the first threshold or below the threshold has indeed an impact. Nonetheless, the thresholds remain around 18%. As $k$ increases, the thresholds for the three setups appear to converge toward the same value and to depend solely on $k$. Table 1 in Appendix B.2 gives the specific values of the thresholds for $k\in\{3,\ldots,10\}$. It is worth noting that the presented setups could be applied to a larger number of groups, although we have limited our analysis to 10 groups.
@@
Once the thresholds for three to ten groups have been identified, the boxplots from the previous section can be reproduced for adjusted scenarios. It is of particular interest to visualize the eigenvalues when one group proportion is at the identified threshold, and to select two or three of the previous mixture proportion scenarios to vary the number of groups with proportions below this threshold. Fig. 7 illustrates this experiment. Although boxplots are plotted, the group centers remain fixed and equal to the values of the first row of the group centers of Table S1 of Supplementary Section S1.
@@
Fig. 7: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ with no within-group variability, with $p=q=k-1$, and 14 different group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$) on the $x$-axis. The values of $k$ vary across panels. In each panel, the second scenario (on the $x$-axis) contains a group at the identified threshold value: approximately 21% when $k=2$, 18% when $k=3$, 14% when $k=5$, and 8% when $k=10$
@@
The results presented in Fig. 7 validate the thresholds of Setup 1 presented in Fig. 6 for $k\in\{2,3,5,10\}$ groups. In each boxplot, the scenario where a group is at the threshold level has an eigenvalue almost equal to one. This is the first eigenvalue, given that these scenarios do not have proportions below the respective threshold. The results are consistent with the logic described in Fig. 5, which states that the number of eigenvalues greater than one is linked to the number of groups with proportion below the threshold. For the scenarios of Fig. 7, when $k=5$ (resp. $k=10$), the number of eigenvalues larger than one is equal to the number of groups with proportions equal to 10% (resp. 5%). However, this relationship needs further investigation in cases involving more than three groups.
@@
In Section 3, Fig. 2 and 3 display the eigenvalues of ICS using $\mathrm{COV}-\mathrm{COV}_{4}$ for various Dirac mixture models described by (5). Fig. 2 illustrates that, if $q=k-1$, these eigenvalues do not depend on the group means, and Proposition 1 in Section 3 confirms the result theoretically. However, as shown in Fig. 3, this result does not hold if $q<k-1$, leading to a more complex analysis of the eigenvalues. In the present section, we examine the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ for model (2) with three Gaussian groups ($k=3$) and $q=1<k-1=2$. This corresponds to the following $p$-dimensional Gaussian mixture:
$\alpha_{1}{\cal N}_{p}(\mathbf{t}_{1},\mathbf{I}_{p})+\alpha_{2}{\cal N}_{p}(\mathbf{t}_{2},\mathbf{I}_{p})+(1-\alpha_{1}-\alpha_{2}){\cal N}_{p}(\mathbf{0}_{p},\mathbf{I}_{p}),$ (6)
with $\mathbf{t}_{1}=(t_{11},0,\ldots,0)^{\top}$, $\mathbf{t}_{2}=(t_{21},0,\ldots,0)^{\top}$, $t_{11}\neq 0$, $t_{21}\neq 0$, and $t_{11}\neq t_{21}$. We can prove the following proposition.
@@
Consider model (6) and let $r_{\alpha_{1},\alpha_{2}}$ be the following polynomial of degree 4:
$r_{\alpha_{1},\alpha_{2}}(x)=\alpha_{1}(-1+7\alpha_{1}-12\alpha_{1}^{2}+6\alpha_{1}^{3})x^{4}+4\alpha_{1}\alpha_{2}(1-6\alpha_{1}+6\alpha_{1}^{2})x^{3}$
$+6\alpha_{1}\alpha_{2}(1-2\alpha_{2}+\alpha_{1}(-2+6\alpha_{2}))x^{2}+4\alpha_{1}\alpha_{2}(1-6\alpha_{2}+6\alpha_{2}^{2})x+\alpha_{2}(-1+7\alpha_{2}-12\alpha_{2}^{2}+6\alpha_{2}^{3}).$
@@
The proof is given in Appendix C. Proposition 2 demonstrates that the behavior of the ICS eigenvalues for a mixture of three Gaussian groups with aligned group centers differs from the two-group case (see Subsection 2.3). For model (6), the condition under which all eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ equal one depends not only on the group proportions but also on the group centers, as determined by the fourth-order (7). This equation has four solutions in the complex plane but, between zero and four solutions in the real space, depending on the group proportions. Using Mathematica (Wolfram Research, Inc., Champaign, IL, 2022) to simplify the computations, we found that for $\alpha_{1}=\alpha_{2}\leq(3-\sqrt{3})/12$ ($\simeq 10.6\%$), or $\alpha_{1}=\alpha_{2}$ equal to $1/3$ or $1/4$, (7) has no real solution. This implies that one and only one eigenvalue of ICS differs from one, and is associated with an eigenvector that spans the FDS (which is only one-dimensional for model (6)). However, we can also identify group proportions for which, given specific group centers, all eigenvalues of ICS equal one, indicating that ICS does not work in all situations. For instance, when $\alpha_{1}=\alpha_{2}=1/6$, (7) has three real solutions, and all the eigenvalues of ICS equal one if $t_{21}/t_{11}$ equal -1, $(2\sqrt{6}+7)/5$ or $(-2\sqrt{6}+7)/5$. While these situations are highly specific, they illustrate that analysing the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ becomes more complex when $q$ is less than $k-1$, and that this phenomenon is already true for $k=3$.